OpenStudy (anonymous):

sqrt(162x^5y^6) Please simplify and provide a simplified version of it

OpenStudy (anonymous):

\[\sqrt{162x^5y^6}=\sqrt{81*2*x^2*x^2*x*y^2*y^2*y^2}=9*x*x*y*y*y \sqrt{2*x}=9x^3y^3\sqrt{2x}\]